A coffee mixture has beans that sell for $0.20 a pound and beans that sell for $0.68. If 120 pounds of beans create a mixture worth $0.54 a pound, how much of each bean is used? Model the scenario then solve it. Then, in two or more sentences explain whether your solution is or is not reasonable.
1
Expert's answer
2020-10-11T18:32:55-0400

Let "x=" the amount of the beans of first type, "y=" the amount of the means of second type. Then


"x+y=120, x\geq0, y\geq0"

The cost of the mixture


"120\cdot0.54=64.8"

"0.20x+0.68y=64.8"

Solve the system


"x+y=120""x+3.4y=324"

Hence


"x+y=120""2.4y=204"

Or


"x=35""y=85"

We mix 85 pounds of beans of high sort (higher cost) and 35 pounds of beans of first sort (lower cost).